# Reliable Decision Support with LLMs: A Framework for Evaluating Consistency in Binary Text Classification Applications

### v1

@@ 1 Introduction
LLM-based annotation has become something of an academic Wild West: the lack of established practices and standards has led to concerns about the quality and validity of research. Researchers have warned that the ostensible simplicity of LLMs can be misleading, as they are prone to bias, misunderstandings, and unreliable results [1, p.1].
LLMs outperform typical human annotators. The evidence is consistent across different types of texts and time periods. It strongly suggests that ChatGPT may already be a superior approach compared to crowd annotations on platforms such as MTurk. At the very least, the findings demonstrate the importance of studying the text-annotation properties and capabilities of LLMs more in depth [2, p.2].
Together, these contrasting perspectives highlight the need to critically examine large language models (LLMs) for text annotation and classification. Although human annotation remains widespread, it poses considerable challenges. It is time-consuming and costly—up to \$5 per annotation and \$50 per hour for annotators [3]—and often suffers from inconsistencies stemming from the intricacies of language and the subjectivity of annotators [4]. Differences in expertise, perspectives, and biases further affect the quality and reliability of human-labeled data [5].
Recent studies suggest that LLMs may address many of these limitations. Evidence shows that LLMs often outperform human annotators in tasks such as relevance assessment, position detection, and topic classification [2, 6, 7, 8, 9, 10]. Tools like GPT-3 and ChatGPT have been shown to reduce annotation costs by 50% to 96%, driven by lower labor needs, faster task completion, and greater scalability [2, 11]. Meanwhile, the rapid adoption of Artificial Intelligence (AI) continues to drive demand for high-quality annotated data. Kniazieva [12] projects that the U.S. data annotation market will grow by 25% annually, reaching \$8.2 billion by 2028. Given this expansion and the persistent challenges of manual annotation, LLMs may substantially disrupt traditional practices.
Despite LLMs’ promise for text annotation, preliminary studies highlight concerns about consistency, reliability, and validity. Evaluating summaries of 100 news articles by 16 LLMs, Stureborg et al. [13, p.1] found LLMs to be “biased evaluators,” exhibiting familiarity bias, uneven rating distributions, and anchoring effects. They also noted low inter-sample agreement and sensitivity to minor prompt variations. Aguda et al. [6] similarly identified a trade-off between LLMs’ efficiency and the precision achieved by experts. Reiss [14] urged caution in using LLMs for classification, emphasizing the need for validity studies. Collectively, the literature highlights the importance of structured methods for effectively integrating LLMs into annotation workflows [1, 15].
@@
Guided by these research questions, we propose a structured framework to support reliable LLM-based annotation. The framework provides guidance on sample size planning, prompt design, model selection criteria, and a methodology to assess reliability and validity [16]. By following these procedures, researchers can evaluate the credibility and applicability of LLM-generated annotations [17, 18]. A financial news sentiment case study demonstrates the utility of the proposed framework. Additionally, we provide open-source code and implementation guidance for broader adoption [17, 18].
@@ 2.1 Large Language Models
Current LLMs are based on the transformer architecture, which efficiently processes large text sequences using self-attention. LLMs can be categorized as (a) encoder-only models like BERT for information extraction [19]; (b) decoder-only models like GPT for fluent text generation [20]; or (c) encoder-decoder models like BART for summarization and translation [21]. LLM outputs are inherently stochastic, with repeated prompts yielding different results.
As LLMs have scaled, they exhibit emergent abilities such as solving math problems and understanding ambiguous language [22]. These capabilities support their classification as foundation models [23], enabling broad application with minimal customization across tasks like text retrieval [24], entity matching [25], sentiment analysis [26], and text classification [27]—the focus of this study. LLMs are increasingly regarded as general-purpose technologies [28], disrupting labor markets [29, 30, 31] and firm performance [32, 33]. However, fundamental uncertainties about their capabilities and failure modes persist [23, 28].
@@ 2.2 LLMs for Text Classification / LLMs-as-a-Judge
Since the release of ChatGPT, LLMs have seen growing use in text classification [20, 34]. Two main factors explain this shift: (1) LLMs can operate in zero- or few-shot settings, reducing reliance on labeled datasets and feature engineering [20, 35]; and (2) NLP methods are foundational across fields such as finance, information systems, and accounting [36, 26, 37]. Comparative studies show that GPT-based models often outperform fine-tuned BERT models, with strong generalization and minimal tuning [15, 36, 38, 39]. However, computational cost, interpretability, and prompt sensitivity remain challenges [1].
Prompt engineering significantly affects classification outcomes [27, 35, 38]. Chain-of-Thought (CoT) prompting, which guides models through reasoning steps, includes zero-shot CoT (“think step by step”) and manual few-shot CoT with examples [40, 41]. Techniques like role-playing can enhance CoT [42]. Manual CoT has been used for fake news detection, job review classification, and grading student writing [42, 28, 33, 43], though no method consistently outperforms others across tasks [1].
A major gap in current research is inconsistent evaluation protocols. Studies often omit replicates, proper sample sizing, or power analysis. For example, Eisfeldt et al. [33] used 100 samples and three replicates without justification. Findings also conflict: Kostina et al. [42, p. 5] found deterministic behavior at temperature 0, contradicting Eisfeldt et al. [33] and Megahed et al. [44]. Verga et al. [45] proposed a diverse panel of models to reduce bias and cost but do not address replication.
@@ 2.3 Intra and Inter-Rater Reliability
Intra-rater reliability captures a single rater’s consistency across repeated assessments, while inter-rater reliability measures agreement across multiple raters [46, 47]. Both are essential for trustworthy measurement, as inconsistency can obscure true values. These concepts are widely used across disciplines such as psychology, healthcare, communication, and NLP [46]. Several statistical measures quantify rater reliability, varying by data type (e.g., categorical vs. continuous) and rater configuration (e.g., two vs. many raters) [48]. Key metrics include:
@@
Cohen’s Kappa ($\kappa$) [49] adjusts for chance agreement between two raters on categorical data:
@@
where $P_{o}$ is observed agreement and $P_{e}$ is expected agreement by chance. $\kappa$ ranges from 1-1 (systematic disagreement) to $1$ (perfect agreement). Landis and Koch [50, p.165] suggest that $\kappa$ values between 0.61 and 0.80 represent “substantial” agreement.
Fleiss’ Kappa [51] extends Cohen’s kappa to multiple raters ($m\geq 3$), assuming the same set of raters rates all items. It provides a chance-corrected agreement measure for nominal data.
Conger’s Kappa [52] also generalizes Cohen’s kappa to multiple raters but allows for varying rater-item assignments.
Brennan-Prediger Coefficient [53] uses an alternative approach to chance correction by assuming uniform category distribution. It is less sensitive to unbalanced marginal distributions than Cohen’s kappa but may still overestimate reliability when the class distributions are unbalanced.
Krippendorff’s Alpha ($\alpha$) [54] is a general-purpose reliability measure that accommodates any number of raters and data types (nominal, ordinal, interval, ratio). It supports missing data and utilizes appropriate distance functions based on the data type. Hayes and Krippendorff [55] advocate $\alpha$ as the standard for content analysis. Törnberg [1] applied $\alpha$ in “prompt stability analysis” for LLMs.
Gwet’s AC1 [56] addresses paradoxes in kappa metrics under unbalanced class distributions. It assumes partial randomness in rater behavior and limits expected chance agreement to 50% per category. While robust to prevalence effects, AC1 may overestimate reliability in balanced datasets [57].
@@ 2.4 Sample Size Calculations
Many foundational reliability measures originated in behavioral sciences, where sample size planning is essential due to the high cost of collecting responses or ratings [46, 48]. Researchers routinely estimate the minimum sample size needed for stable reliability estimates, particularly when assessing internal consistency or rater agreement. In contrast, LLM-based classification tasks often involve massive corpora. For example, Sutherland et al. [58] analyzed over 530,000 CPA profiles, illustrating the scale of LLM-assisted annotation. Downsampling enables cost-effective, statistically grounded evaluations of model reliability and validity. Using a smaller, representative sample avoids wasting resources on unreliable models. Once a model demonstrates sufficient consistency and validity, it can be confidently applied to the full dataset. Our approach supports this workflow by providing statistical tools to estimate sample sizes needed for robust reliability assessment.
@@ 3 Our Proposed Framework
Our framework for evaluating LLM-based text classification includes four phases (Figure 1): planning, data collection, reliability analysis, and validity analysis. The planning phase determines the minimum sample size, selects LLMs, and designs prompts. Data collection involves curating examples, implementing code infrastructure, and running the annotation experiment. Reliability analysis assesses annotation consistency across runs, while validity analysis measures accuracy against benchmark or external criterion labels.
@@ 3.1 Planning Phase
Evaluating the reliability of binary text annotation requires careful planning due to the complexity of the task. We recommend computing the following complementary reliability metrics: Conger’s Kappa, Fleiss’ Kappa, Gwet’s AC1, Brennan-Prediger Coefficient, and Krippendorff’s Alpha. Using multiple metrics ensures a more comprehensive evaluation, as no single measure fully captures the nuances of reliability in binary annotation. These metrics can be efficiently computed using open-source tools such as the irrCAC package [59], enabling reproducible and systematic assessment for similar tasks.
@@ 3.1.1 Sample Size Determination
To determine the minimum required examples, we adapted Gwet’s [48] psychometric approach. The examples to be annotated are treated as analogous to “subjects,” while the “raters” represent either multiple LLMs (inter-reliability) or multiple replicates of a single LLM’s outputs (intra-reliability).
We selected a target margin of error ($E_{0}=0.10$) and a confidence level of ($1-\alpha=0.90$), standard values in psychometric research. From the psychometric literature, sample size estimation techniques are only presented for three reliability metrics: Percent Agreement, Gwet’s AC1, and Brennan-Prediger Coefficient [48]. For each metric, the sample size can be calculated as:
@@
where $z_{\alpha}$ is the critical value corresponding to the confidence level ($\approx 1.645$ for 90%), $C$ is a parameter that depends on the reliability metric, number of raters/replicates ($r$), and the number of categories ($q$), and $E_{0}$ is the desired margin of error. This ensures the sample size is large enough to achieve the desired level of precision by controlling the standard error of the agreement coefficient. The $C$ values can be computed as $C=\frac{1}{A}$, where $A$ is the agreement value for a given metric. Values of $A$ can be obtained from the agreement tables provided in Gwet’s handbook [48], using the number of categories ($q$) and the number of replicates/raters ($r$). The tables show that $A$ increases with $q$ and $r$. For binary classification, $q=2$ since the output is binary. Conversely, $r$ should be set to the number of planned replicates per model in the case of intra-rater reliability. By inserting the values for $C$, $E_{0}$, and $z_{\alpha}$, we obtain:
@@ 3.1.2 Selecting Candidate LLMs
Selecting candidate LLMs requires careful consideration of several factors. First, data sensitivity may necessitate locally deployed models rather than API-based services. Second, computational capacity and associated costs can significantly impact the feasibility of large-scale annotation. Third, expected model performance should be evaluated using established benchmarks and leaderboards [60]. Together, these considerations help identify models that align with task needs and practical constraints.
@@ 3.1.3 Prompt Design and Engineering
Prompt design for text annotation includes two key components [1]: a detailed codebook and effective prompt engineering. The codebook should define categories and outline how to handle edge cases [61, 1]. Prompt engineering should use techniques like manual CoT to elicit step-by-step reasoning [62, 63]. Both the codebook and prompts should be iteratively refined through testing to ensure the outputs meet expectations.
@@ 3.2.1 Curating Representative Examples for Annotation
The first step is to curate a dataset with at least $n_{\text{final}}$ examples, prioritizing diverse and edge cases relevant to the task [10]. If no benchmark exists, a stratified random sample should be drawn from the full dataset. The curated set should be class-balanced and include both the text to be annotated and, when available, benchmark or external criterion labels.
@@ 3.2.2 Setting Up the LLM-based Classification System
Third, we use the irrCAC package [59] to compute reliability metrics and agreement scores. This component must be designed to process invalid responses outside the prescribed annotation categories (e.g., “neutral” when only “positive” or “negative” options are specified).

### v2

@@ 1 Introduction
LLM-based annotation has become something of an academic Wild West: the lack of established practices and standards has led to concerns about the quality and validity of research. Researchers have warned that the ostensible simplicity of LLMs can be misleading, as they are prone to bias, misunderstandings, and unreliable results [tornberg2024best, p.1].
LLMs outperform typical human annotators. The evidence is consistent across different types of texts and time periods. It strongly suggests that ChatGPT may already be a superior approach compared to crowd annotations on platforms such as MTurk. At the very least, the findings demonstrate the importance of studying the text-annotation properties and capabilities of LLMs more in depth [gilardi2023chatgpt, p.2].
Together, these contrasting perspectives highlight the need to critically examine large language models (LLMs) for text annotation and classification. Although human annotation remains widespread, it poses considerable challenges. It is time-consuming and costly—up to \$5 per annotation and \$50 per hour for annotators [labelers2025explore]—and often suffers from inconsistencies stemming from the intricacies of language and the subjectivity of annotators [wu2024modelling]. Differences in expertise, perspectives, and biases further affect the quality and reliability of human-labeled data [davani2022dealing].
Recent studies suggest that LLMs may address many of these limitations. Evidence shows that LLMs often outperform human annotators in tasks such as relevance assessment, position detection, and topic classification [gilardi2023chatgpt, aguda2024large, chiang2023can, huang2023chatgpt, thapa2023humans, tornberg2023chatgpt]. Tools like GPT-3 and ChatGPT have been shown to reduce annotation costs by 50% to 96%, driven by lower labor needs, faster task completion, and greater scalability [gilardi2023chatgpt, wang2021want]. Meanwhile, the rapid adoption of Artificial Intelligence (AI) continues to drive demand for high-quality annotated data. Kniazieva [kniazieva2024latest] projects that the U.S. data annotation market will grow by 25% annually, reaching \$8.2 billion by 2028. Given this expansion and the persistent challenges of manual annotation, LLMs may substantially disrupt traditional practices.
Despite LLMs’ promise for text annotation, preliminary studies highlight concerns about consistency, reliability, and validity. Evaluating summaries of 100 news articles by 16 LLMs, Stureborg et al. [stureborg2024large, p.1] found LLMs to be “biased evaluators,” exhibiting familiarity bias, uneven rating distributions, and anchoring effects. They also noted low inter-sample agreement and sensitivity to minor prompt variations. Aguda et al. [aguda2024large] similarly identified a trade-off between LLMs’ efficiency and the precision achieved by experts. Reiss [reiss2023testing] urged caution in using LLMs for classification, emphasizing the need for validity studies. Collectively, the literature highlights the importance of structured methods for effectively integrating LLMs into annotation workflows [tornberg2024best, pangakis2023automated].
@@
Guided by these research questions, we propose a structured framework to support reliable LLM-based annotation. The framework provides guidance on sample size planning, prompt design, model selection criteria, and a methodology to assess reliability and validity [alizadeh2024open]. By following these procedures, researchers can evaluate the credibility and applicability of LLM-generated annotations [liu2024poliprompt, nie2024survey]. A financial news sentiment case study demonstrates the utility of the proposed framework. Additionally, we provide open-source code and implementation guidance for broader adoption [liu2024poliprompt, nie2024survey].
@@ 2.1 Large Language Models
Current LLMs are based on the transformer architecture, which efficiently processes large text sequences using self-attention. LLMs can be categorized as (a) encoder-only models like BERT for information extraction [devlin2019bert]; (b) decoder-only models like GPT for fluent text generation [brown2020language]; or (c) encoder-decoder models like BART for summarization and translation [lewis2020bart]. LLM outputs are inherently stochastic, with repeated prompts yielding different results.
As LLMs have scaled, they exhibit emergent abilities such as solving math problems and understanding ambiguous language [wei2022characterizing]. These capabilities support their classification as foundation models [bommasani2022opportunities], enabling broad application with minimal customization across tasks like text retrieval [shi2025know], entity matching [babaian2024entity], sentiment analysis [mousavi2024resilience], and text classification [sun2023text]—the focus of this study. LLMs are increasingly regarded as general-purpose technologies [eloundou2024gpts], disrupting labor markets [wef2025future, handa2024economic, demirci2025ai] and firm performance [dell2023navigating, eisfeldt2023generative, reusens2025llm]. However, fundamental uncertainties about their capabilities and failure modes persist [salazar2025contribution, bommasani2022opportunities, eloundou2024gpts].
@@ 2.2 LLMs for Text Classification / LLMs-as-a-Judge
Since the release of ChatGPT, LLMs have seen growing use in text classification [brown2020language, ouyang2022training]. Two main factors explain this shift: (1) LLMs can operate in zero- or few-shot settings, reducing reliance on labeled datasets and feature engineering [brown2020language, chae2023large]; and (2) NLP methods are foundational across fields such as finance, information systems, and accounting [maibaum2024selecting, mousavi2024resilience, shao2025revisiting]. Comparative studies show that GPT-based models often outperform fine-tuned BERT models, with strong generalization and minimal tuning [pangakis2023automated, maibaum2024selecting, peeters2023entity, liu2023pre]. However, computational cost, interpretability, and prompt sensitivity remain challenges [tornberg2024best].
Prompt engineering significantly affects classification outcomes [sun2023text, chae2023large, peeters2023entity]. Chain-of-Thought (CoT) prompting, which guides models through reasoning steps, includes zero-shot CoT (“think step by step”) and manual few-shot CoT with examples [wei2022chain, Weng2023-az]. Techniques like role-playing can enhance CoT [kostina2025large]. Manual CoT has been used for fake news detection, job review classification, and grading student writing [kostina2025large, eloundou2024gpts, eisfeldt2023generative, impey2025using], though no method consistently outperforms others across tasks [tornberg2024best].
A major gap in current research is inconsistent evaluation protocols. Studies often omit replicates, proper sample sizing, or power analysis. For example, Eisfeldt et al. [eisfeldt2023generative] used 100 samples and three replicates without justification. Findings also conflict: Kostina et al. [kostina2025large, p. 5] found deterministic behavior at temperature 0, contradicting Eisfeldt et al. [eisfeldt2023generative] and Megahed et al. [megahed2024introducing]. Verga et al. [verga2024replacing] proposed a diverse panel of models to reduce bias and cost but do not address replication.
@@ 2.3 Intra and Inter-Rater Reliability
Intra-rater reliability captures a single rater’s consistency across repeated assessments, while inter-rater reliability measures agreement across multiple raters [krippendorff2019content, shrout1979intraclass]. Both are essential for trustworthy measurement, as inconsistency can obscure true values. These concepts are widely used across disciplines such as psychology, healthcare, communication, and NLP [krippendorff2019content]. Several statistical measures quantify rater reliability, varying by data type (e.g., categorical vs. continuous) and rater configuration (e.g., two vs. many raters) [gwet2021handbook]. Key metrics include:
@@
Cohen’s Kappa ($\kappa$) [cohen1960coefficient] adjusts for chance agreement between two raters on categorical data:
@@
where $P_{o}$ is observed agreement and $P_{e}$ is expected agreement by chance. $\kappa$ ranges from 1-1 (systematic disagreement) to $1$ (perfect agreement). Landis and Koch [landis1977measurement, p.165] suggest that $\kappa$ values between 0.61 and 0.80 represent “substantial” agreement.
Fleiss’ Kappa [fleiss1971measuring] extends Cohen’s kappa to multiple raters ($m\geq 3$), assuming the same set of raters rates all items. It provides a chance-corrected agreement measure for nominal data.
Conger’s Kappa [conger1980integration] also generalizes Cohen’s kappa to multiple raters but allows for varying rater-item assignments.
Brennan-Prediger Coefficient [brennan1981coefficient] uses an alternative approach to chance correction by assuming uniform category distribution. It is less sensitive to unbalanced marginal distributions than Cohen’s kappa but may still overestimate reliability when the class distributions are unbalanced.
Krippendorff’s Alpha ($\alpha$) [krippendorff1970bivariate] is a general-purpose reliability measure that accommodates any number of raters and data types (nominal, ordinal, interval, ratio). It supports missing data and utilizes appropriate distance functions based on the data type. Hayes and Krippendorff [hayes2007answering] advocate $\alpha$ as the standard for content analysis. Törnberg [tornberg2024best] applied $\alpha$ in “prompt stability analysis” for LLMs.
Gwet’s AC1 [gwet2002inter] addresses paradoxes in kappa metrics under unbalanced class distributions. It assumes partial randomness in rater behavior and limits expected chance agreement to 50% per category. While robust to prevalence effects, AC1 may overestimate reliability in balanced datasets [gwet2008computing].
@@ 2.4 Sample Size Calculations
Many foundational reliability measures originated in behavioral sciences, where sample size planning is essential due to the high cost of collecting responses or ratings [krippendorff2019content, gwet2021handbook]. Researchers routinely estimate the minimum sample size needed for stable reliability estimates, particularly when assessing internal consistency or rater agreement. In contrast, LLM-based classification tasks often involve massive corpora. For example, Sutherland et al. [sutherland2024occupational] analyzed over 530,000 CPA profiles, illustrating the scale of LLM-assisted annotation. Downsampling enables cost-effective, statistically grounded evaluations of model reliability and validity. Using a smaller, representative sample avoids wasting resources on unreliable models. Once a model demonstrates sufficient consistency and validity, it can be confidently applied to the full dataset. Our approach supports this workflow by providing statistical tools to estimate sample sizes needed for robust reliability assessment.
@@ 3 Our Proposed Framework
Our framework for evaluating LLM-based text classification includes four phases (Figure 1): planning, data collection, reliability analysis, and validity analysis. The planning phase determines the minimum sample size, selects LLMs, and designs prompts. Data collection involves curating examples, implementing code infrastructure, and running the annotation experiment. Reliability analysis assesses annotation consistency across runs, while validity analysis measures accuracy against benchmark or external criterion labels. The framework is designed to be domain-agnostic and can be applied to any binary classification task.
@@ 3.1 Planning Phase
Evaluating the reliability of binary text annotation requires careful planning due to the complexity of the task. We recommend computing the following complementary reliability metrics: Conger’s Kappa, Fleiss’ Kappa, Gwet’s AC1, Brennan-Prediger Coefficient, and Krippendorff’s Alpha. Using multiple metrics ensures a more comprehensive evaluation, as no single measure fully captures the nuances of reliability in binary annotation. These metrics can be efficiently computed using open-source tools such as the irrCAC package [gwet2019irrCAC], enabling reproducible and systematic assessment for similar tasks.
@@ 3.1.1 Sample Size Determination
To determine the minimum required examples, we adapted Gwet’s [gwet2021handbook] psychometric approach. The examples to be annotated are treated as analogous to “subjects,” while the “raters” represent either multiple LLMs (inter-reliability) or multiple replicates of a single LLM’s outputs (intra-reliability).
We selected a target margin of error ($E_{0}=0.10$) and a confidence level of ($1-\alpha=0.90$), standard values in psychometric research. From the psychometric literature, sample size estimation techniques are only presented for three reliability metrics: Percent Agreement, Gwet’s AC1, and Brennan-Prediger Coefficient [gwet2021handbook]. For each metric, the sample size can be calculated as:
@@
where $z_{\alpha}$ is the critical value corresponding to the confidence level ($\approx 1.645$ for 90%), $C$ is a parameter that depends on the reliability metric, number of raters/replicates ($r$), and the number of categories ($q$), and $E_{0}$ is the desired margin of error. This ensures the sample size is large enough to achieve the desired level of precision by controlling the standard error of the agreement coefficient. The $C$ values can be computed as $C=\frac{1}{A}$, where $A$ is the agreement value for a given metric. Values of $A$ can be obtained from the agreement tables provided in Gwet’s handbook [gwet2021handbook], using the number of categories ($q$) and the number of replicates/raters ($r$). The tables show that $A$ increases with $q$ and $r$. For binary classification, $q=2$ since the output is binary. Conversely, $r$ should be set to the number of planned replicates per model in the case of intra-rater reliability. By inserting the values for $C$, $E_{0}$, and $z_{\alpha}$, we obtain:
@@ 3.1.2 Selecting Candidate LLMs
Selecting candidate LLMs requires careful consideration of several factors. First, data sensitivity may necessitate locally deployed models rather than API-based services. Second, computational capacity and associated costs can significantly impact the feasibility of large-scale annotation. Third, expected model performance should be evaluated using established benchmarks and leaderboards [chiang2024chatbot]. Together, these considerations help identify models that align with task needs and practical constraints.
@@ 3.1.3 Prompt Design and Engineering
Prompt design for text annotation includes two key components [tornberg2024best]: a detailed codebook and effective prompt engineering. The codebook should define categories and outline how to handle edge cases [glaser1999discovery, tornberg2024best]. Prompt engineering should use techniques like manual CoT to elicit step-by-step reasoning [amatriain2024prompt, minaee2024large]. Both the codebook and prompts should be iteratively refined through testing to ensure the outputs meet expectations.
@@ 3.2.1 Curating Representative Examples for Annotation
The first step is to curate a dataset with at least $n_{\text{final}}$ examples, prioritizing diverse and edge cases relevant to the task [tornberg2023chatgpt]. If no benchmark exists, a stratified random sample should be drawn from the full dataset. The curated set should be class-balanced and include both the text to be annotated and, when available, benchmark or external criterion labels.
@@ 3.2.2 Setting Up the LLM-based Classification System
Third, we use the irrCAC package [gwet2019irrCAC] to compute reliability metrics and agreement scores. This component must be designed to process invalid responses outside the prescribed annotation categories (e.g., “neutral” when only “positive” or “negative” options are specified).